OpenStudy (anonymous):

prove: 2cosx+2cos(square)x over sin2x = sinx over 1-cosx ???

OpenStudy (vishal_kothari):

can you clarify it more....

OpenStudy (lollylau):

"confused"

OpenStudy (turingtest):

\[\frac{2\cos x+2\cos^2x}{\sin(2x)}=\frac{\sin x}{1-\cos^2x}\]this or no?

OpenStudy (anonymous):

YES!!!! :) thanks.

OpenStudy (turingtest):

no wait, you have\[\frac{2\cos x+2\cos^2x}{\sin(2x)}=\frac{\sin x}{1-\cos x}\] which is it? you need to be sure, or else we can't help.

OpenStudy (turingtest):

(I took the square off the cosine on the right)

OpenStudy (anonymous):

Oo! Second one. :)

OpenStudy (turingtest):

oh I got it...

OpenStudy (anonymous):

\[\frac{2\cos x (1 + \cos x)}{\sin (2x)} = \frac{\sin x}{1-\cos x}\]

OpenStudy (anonymous):

divide 1+ cos x on both sides

OpenStudy (anonymous):

1 - cos^2 x = sin^2 x

OpenStudy (anonymous):

then you get sin (2x) = 2 cos x sin x

OpenStudy (anonymous):

ohh yay thank you guys!!!! :)

OpenStudy (anonymous):

I never know how to start trig identity questions.. is there any stragedy to solve it? :P

OpenStudy (anonymous):

always convert tangent in terms of cosine and sine

OpenStudy (turingtest):

things like that (2x) have to be on both sides or neither. Definitely need to take care of that in my opinion. That said I did it slightly differently than moneybird, so there are different ways. I can post it but it would take a sec..

OpenStudy (turingtest):

I mean the (2x) in the sine in the denominator...

OpenStudy (anonymous):

I would appreciate your solution! :)

OpenStudy (turingtest):

I think I cut moneybird off, he wasn't done yet... sorry moneybird

OpenStudy (anonymous):

(1 - cos x) (1 + cos x) = 1 - cos^2 x

OpenStudy (mertsj):

If this is an identity that is supposed to be proven, you cannot treat it like an equation where you do the same thing to both sides, moneybird.

OpenStudy (anonymous):

moneybird, I really appreciate your help! I am still confused what has happened.. so [2cos ^2x becomes (1+cosx)]?

OpenStudy (turingtest):

\[\frac{2\cos x+2\cos^2x}{\sin(2x)}=\frac{\sin x}{1-\cos x}\]like I said take care of the sin(2x) with the formula\[\sin(2x)=2\sin x\cos x\]so we get\[\frac{2\cos x+2\cos^2x}{2\sin x\cos x}=\frac{\sin x}{1-\cos x}\]cancel out a 2 and a cosine on the left\[\frac{1+\cos x}{\sin x}=\frac{\sin x}{1-\cos x}\]Now we want to see if we can get something familiar out of this, like an identity we know. It may take practice just to see the next moves coming: multiply by the denominator of each side\[(1-\cos x)(1+\cos x)=\sin^2x\]\[1-\cos^2x=\sin^2x\]\[1=\sin^2x+\cos^2x\]which is always true.

OpenStudy (anonymous):

cancel out a 2 and a cosine on the left?

OpenStudy (mertsj):